Mass of crucible and cover (g) 31.507 Mass of crucible, cover, and alum before heating (g) Mass of crucible, cover, and alum after 1st heating (g) 1 Mass of crucible, cover, and alum after 2nd heating (g) | Mass of crucible, cover, and alum after 3rd heating (g) Mass of anhydrous alum (g) 3 2, C 2
DATA ANALYSIS Based on your experimental data, calculate the percent water in your sample. 2. Based on your experimental data calculated the value of n in the chemical formula nHO using a periodic 3. Calculate the theoretical percentage of SO, in pure KAKSO table for atomic weights and the correct literature value for n. 4. Is your sample alum? Use the results of the three tests to support your answer. Discuss the accuracy of your tests and possible sources of experimental error 5. If the melting temperature test was the only test that you conducted, how confident would you be in the identification of your sample? Explain. Classity the tests you performed as either physi Melting point: Water of hydration Test for sulfate: Test for aluminum: property . qualitative or quantitati
Answer #1

Mass of anhydrous alum = ( Mass of crucible, cover and alum after 3rd heating) - (Mass of crucible, cover) = 32.421 - 31.507 = 0.914 g

1. Mass of water = Mass of alum - mass of anhydrous alum

Mass of alum = Mass of crucible, cover and alum after before heating) - (Mass of crucible, cover) = 33.287 - 31.507 = 1.780 g

Mass of water = 1.780 - 0.914 = 0.866 g

Mass % of water in sample = (Mass of water /Mass of sample)*100% = (0.866 g/ 1.780)*100% = 48.6%

2. Molar mass of KAl(SO4)2 = Mass of K + Mass of Al + (2* mass of S) + (8*Mass of O) = 39.0983 + 26.9815 + (2*32.065) + (8*15.9994) = 258.205 g/mol

Molar mass of water = 18.02 g/mol

Moles of KAl(SO4)2 = Mass/ molar mass = 0.914/(258.205 g/mol) = 0.00354 mol

Moles of water = 0.866 / (18.02 g/mol) = 0.0481 mol

Mole ratio of KAl(SO4)2 and H2O

0.00354 mol KAl(SO4)2 : 0.0481 mol H2O

0.00354 mol KAl(SO4)2/ 0.00354 mol : 0.0481 mol H2O/ 0.00354 mol

1 : 14

So, n = 14

3.Theoretical value of alum is KAl(SO4)2. 12 H2O

1 mol of alum contains 2 mol of SO42-

Mass of SO42- = moles * molar mass = 2 *96.0626 = 192.1252g

Theoretical percentage of SO42- = (192.1252 g / 474.3884) *100% = 40.5%